# Multiple ergodic averages along functions from a Hardy field: convergence, recurrence and combinatorial applications

### v2

@@ 1.1.   Combinatorial results
1.2 was later improved in [BLL08] to give an “if and only if” condition. A finite collection of polynomials $q_{1},\ldots,q_{k}\in\mathbb{Z}[t]$ is called jointly intersective if for all $m\in\mathbb{N}$ there is $n\in\mathbb{N}$ such that $q_{1}(n)\equiv\ldots\equiv q_{k}(n)\equiv{0}\bmod{m}$.
@@
In view of the above results, one is led to inquire whether similar results hold for more general classes of, say, eventually monotone sequences that do not grow too fast [^2] We remark in passing that for any sequence of exponential growth $q(n)$ there exists a set $E\subset\mathbb{N}$ with $\overline{\mathrm{d}}(E)>0$ which contains no pair of the form $\{x,x+q(n)\}$ with $x,n\in\mathbb{N}$ (cf. [Bei+08, Corollary 4.18]).. A natural class of sequences to consider are those arising from Hardy fields.
@@
Let $\mathsf{G}$ denote the ring (under pointwise addition and multiplication) of germs at infinity [^3] A germ at infinity is any equivalence class of real-valued functions in one real variable under the equivalence relationship $(f\equiv g)\Leftrightarrow\big{(}\exists t_{0}>0~{}\text{such that}~{}f(t)=g(t)~{}\text{for all}~{}t\in[t_{0},\infty)\big{)}$. of real valued functions defined on a half-line $[s,\infty)$ for some $s\in\mathbb{R}$. Any subfield of $\mathsf{G}$ that is closed under differentiation is called a Hardy field.
By abuse of language, we say that a function $f\colon[s,\infty)\to\mathbb{R}$ belongs to some Hardy field $\mathcal{H}$, and write $f\in\mathcal{H}$, if its germ at infinity belongs to $\mathcal{H}$. For convenience, we assume all Hardy fields considered in this paper are shift-invariant, i.e., if $f\in\mathcal{H}$ then $f_{u}\in\mathcal{H}$ for all $u\in\mathbb{R}$, where $f_{u}(t)=f(t+u)$. A classical example of a (shift-invariant) Hardy field is the class of logarithmico-exponential functions introduced by Hardy in [Har12, Har71]. It consists of all (germs of) real-valued functions that can be built from real polynomials, the logarithmic function $\log(t)$, and the exponential function $\exp(t)$ using the standard arithmetical operations $+$, --, $\cdot$, $\div$ and the operation of composition. Examples of logarithmico-exponential functions are ${p(t)}/{q(t)}$ for $p(t),q(t)\in\mathbb{R}[t]$, $t^{c}$ for $c\in\mathbb{R}$, ${t}/{\log(t)}$ and $e^{\sqrt{t}}$, as well as any products or linear combinations of the above. Other Hardy fields contain even more exotic functions, such as $e^{\sqrt{\log t}}\Gamma(t)$ and $t^{c}\zeta(t)$ for any $c\in\mathbb{R}$, where $\zeta$ is the Riemann zeta function and $\Gamma$ is the usual gamma function (cf. [Bos84]). For more information on Hardy fields we refer the reader to [Bos81, Bos94, Fra09].
@@
Another variant of Szemerédi’s Theorem, which was recently obtained by the authors in [BMR20], stands in general position to 1.5 and reveals a new phenomenon pertaining to multiple recurrence along a wide family of non-polynomial (and not necessarily Hardy) functions. The following is one of the combinatorial corollaries of the main result in [BMR20].
@@
Both Theorems 1.5 and 1.6 leave some space for further inquiry. For instance, one would like to know if 1.6 has a version for several functions from $\mathcal{H}$ which deals with the patterns
@@
As for 1.5, a natural generalization is addressed by the following conjecture of Frantzikinakis. Given a finite collection of functions $f_{1},\ldots,f_{k}$, define
$\mathrm{span}(f_{1},\ldots,f_{k})\coloneqq\big{\{}c_{1}f_{1}(t)+\ldots+c_{k}f_{k}(t):(c_{1},\ldots,c_{k})\in\mathbb{R}^{k}\big{\}}.$
@@
$\mathrm{span}^{*}(f_{1},\ldots,f_{k})\coloneqq\big{\{}c_{1}f_{1}(t)+\ldots+c_{k}f_{k}(t):(c_{1},\ldots,c_{k})\in\mathbb{R}^{k}\backslash\{0\}\big{\}}.$
@@
In the statements of Theorems 1.5 and 1.6 and 1.7, it is possible to replace the floor function $\lfloor.\rfloor\colon\mathbb{R}\to\mathbb{Z}$ with other rounding functions, such as the ceiling function $\lceil\cdot\rceil\colon\mathbb{R}\to\mathbb{Z}$, or the rounding to the closest integer function $[.]\colon\mathbb{R}\to\mathbb{Z}$. Indeed, since $\lceil x\rceil=-\lfloor-x\rfloor$ and $[x]=\lfloor x+0.5\rfloor$, replacing $\lfloor$.$\rfloor$ . ⌋ with either $\lceil$.$\rceil$ . ⌉ or [.][.][ . ] actually yields equivalent formulations of those statements.
The following theorem is the main combinatorial result of this paper. It contains Theorems 1.1, 1.2, 1.3, 1.5, 1.6 as special cases and confirms 1.7. We denote by $\mathrm{poly}(f_{1},\ldots,f_{k})$ the set of all real polynomials that can be “generated” using linear combinations of the functions $f_{1},\ldots,f_{k}$. More precisely,
@@
Condition 1 in A will still hold if all the $f_{i}$ are shifted. Together with 1.8, this observation implies that if we are in case 1 of A then the conclusion of A remains true, even if the closest integer function []delimited-[][$\cdot$][ ⋅ ] is replaced by either $\lfloor$.$\rfloor$ . ⌋ or $\lceil$.$\rceil$ . ⌉ (or indeed with any other rounding function of the form $\lfloor x+c\rfloor$, $c\in[0,1)$). This observation, however, does not apply to condition 2 – see 1.14 below.
Besides resolving 1.7, A also implies numerous new results. Corollaries A1, A2, A3, and A4 below comprise a selection of such results that we consider to be of particular interest.
The following rather special case of A already implies Theorems 1.1, 1.2, 1.5, and 1.6.
@@
where $c_{1},\ldots,c_{k}$ are all positive integers, follows from 1.2, whereas the case when $c_{1},\ldots,c_{k}$ are all positive non-integers follows from 1.5. The next corollary of A deals with the previously unknown case where the constants $c_{1},\ldots,c_{k}$ are a mix of integers and non-integers.
@@
For any $c_{1},\ldots,c_{k}>0$ and any set $E\subset\mathbb{N}$ of positive upper density there exist $a,n\in\mathbb{N}$ such that $\{a,\,a+[n^{c_{1}}],\ldots,a+[n^{c_{k}}]\}\subset E$. The same is true with [.][.][ . ] replaced by either $\lfloor$.$\rfloor$ . ⌋ or $\lceil$.$\rceil$ . ⌉.
@@
We remark that Corollaries A1 and A3 are not true if [.][.][ . ] is replaced by either $\lfloor$.$\rfloor$ . ⌋ or $\lceil$.$\rceil$ . ⌉ (see 1.14 below).
While Corollary A3 contains 1.2 as a special case, it does not encompass polynomials which have a non-zero constant term. The following theorem shows that the conclusion of Corollary A3 holds for significantly more general families ${\mathcal{K}}$. In particular, in conjunction with A, it implies both 1.3 and 1.5.
@@
Let $q\in\mathbb{Z}[t]$ be an intersective polynomial, let $\mathcal{H}$ be a Hardy field and let ${\mathcal{L}}\subset\mathcal{H}$ be a family of functions such that any $f\in{\mathcal{L}}$ satisfies $t^{k-1}\prec f(t)\prec t^{k}$ for some $k\in\mathbb{N}$, and any distinct $f,g\in{\mathcal{L}}$ have different growth. Let ${\mathcal{K}}({\mathcal{L}},q)$ be the linear span over $\mathbb{R}$ of ${\mathcal{L}}$ and $q\mathbb{R}[t]$. Then any tuple $f_{1},\dots,f_{k}$ from ${\mathcal{K}}({\mathcal{L}},q)$ satisfies Condition 2 of A.
@@
-$\nabla\text{-}\mathrm{span}(f_{1},\ldots,f_{k})$$\coloneqq\big{\{}c_{1}f_{1}^{(m_{1})}(t)+\ldots+c_{k}f_{k}^{(m_{k})}(t):c_{1},\ldots,c_{k}\in\mathbb{R},\,m_{1},\ldots,m_{k}\in\mathbb{N}\cup\{0\}\big{\}}.$
A allows us to derive a corollary which extends 1.6 from a single function $f$ to multiple functions $f_{1},\ldots,f_{k}$.
@@
The same is true with [.][.][ . ] replaced by either $\lfloor$.$\rfloor$ . ⌋ or $\lceil$.$\rceil$ . ⌉.
The assumptions of Corollary A4 are satisfied, for instance, if all the $f_{i}$ are a linear combinations of powers $t^{c}$ with non-integer exponents $c>0$. On the other hand, if some $f\in\mathrm{span}^{*}(f_{1},\dots,f_{k})$ is a polynomial, then the conclusion of Corollary A4 fails. In Example 1.20 below we show that the conclusion may fail even when $\mathrm{poly}(f_{1},\dots,f_{k})=\emptyset$.
@@ 1.2.   Ergodic results
In light of (1.2), it is natural to tackle A by studying the behaviour of Cesàro averages of multicorrelation expressions of the form
@@
It is important to mention that given a finite collection of functions $f_{1},\ldots,f_{k}\in\mathcal{H}$ of polynomial growth from a Hardy field $\mathcal{H}$, not every $W\in\mathcal{H}$ will satisfy property 1. For example $W(t)=t$ is not compatible with $f(t)=\log t$. On the other hand, there always exists a compatible $W\in\mathcal{H}$ (see 2.2). Roughly speaking, $W$ can be chosen as any function which tends to $\infty$ and grows “slower” than any unbounded function in
@@
Let $f_{1},\ldots,f_{k}$ be functions of polynomial growth from a Hardy field $\mathcal{H}$ and let $W\in\mathcal{H}$ be any function with $1\prec W(t)\ll t$ and such that $f_{1},\ldots,f_{k}$ satisfy Property 1. Let $(X,\mathcal{B},\mu,T)$ be an invertible measure preserving system.
  1. [label=(),ref=(),leftmargin=*]
  2. 1.
@@
  3. 2.
    If $f_{1},\ldots,f_{k}$ satisfy condition 1 of A then for any $h_{1},\ldots,h_{k}\in L^{\infty}(X)$
[…]
@@
  4. 3.
    If $f_{1},\ldots,f_{k}$ satisfy condition 2 of A then for any $A\in\mathcal{B}$ with $\mu(A)>0$ we have
[…]
As a first corollary to B we obtain convergence of multiple ergodic averages with Cesàro weights along a rather large class of functions from a Hardy field. Special cases of this corollary were previously obtained in [HK05a, HK05, Lei05, BH09, Fra10, Kou18].
@@
exists in $L^{2}$. The same is true with [.][.][ . ] replaced by either $\lfloor$.$\rfloor$ . ⌋ or $\lceil$.$\rceil$ . ⌉.
B also provides additional information on the following result, which was originally conjectured by Frantzikinakis ([Fra16, Problem 23]; cf. also [Fra15, Problem 1]) and recently proved by Tsinas [Tsi23].
@@
Part 2 of B implies 1.12 under the additional assumption that the functions $f_{1},\dots,f_{k}$ satisfy Property 1 with $W(t)=t$, which is slightly stronger than the assumption that $|f(t)-q(t)|/\log t\to\infty$ for every $q\in\mathbb{Z}[t]$ and $f\in\mathrm{span}^{*}(f_{1},\ldots,f_{k})$. However, B gives more than that. Indeed, part 2 of B provides a confirmation of a variant of 1.12 where the condition $|f(t)-q(t)|/\log t\to\infty$ is weakened to $|f(t)-q(t)|\to\infty$, but at the price of replacing Cesàro averages by weighted averages (where one can choose any weight for which Property 1 is satisfied, cf. 2.2).
@@
Let $f_{1},\ldots,f_{k}$ be functions of polynomial growth from a Hardy field $\mathcal{H}$ and let $W\in\mathcal{H}$ be any function with $1\prec W(t)\ll t$ and such that $f_{1},\ldots,f_{k}$ satisfy Property 1. Suppose that for all $q\in\mathbb{Z}[t]$ and $f\in\mathrm{span}^{*}(f_{1},\ldots,f_{k})$ we have $\lim_{t\to\infty}|f(t)-q(t)|=\infty$. Then for any invertible ergodic measure preserving system $(X,\mathcal{B},\mu,T)$ and any $h_{1},\dots,h_{k}\in L^{\infty}(X)$
@@
The same is true when [.][.][ . ] is replaced by either $\lfloor$.$\rfloor$ . ⌋ or $\lceil$.$\rceil$ . ⌉.
@@
Finally, we formulate two corollaries of B which imply A (via Furstenberg’s correspondence principle). We believe that these two results are also of independent interest.
@@
Let $f_{1},\ldots,f_{k}$ be functions of polynomial growth from a Hardy field $\mathcal{H}$ satisfying condition 1 of A. Then for any invertible measure preserving system $(X,\mathcal{B},\mu,T)$ and any $A\in\mathcal{B}$,
@@
The same is true when [.][.][ . ] is replaced by either $\lfloor$.$\rfloor$ . ⌋ or $\lceil$.$\rceil$ . ⌉.
@@
Let $f_{1},\ldots,f_{k}$ be functions of polynomial growth from a Hardy field $\mathcal{H}$ satisfying condition 2 of A. Then for any invertible measure preserving system $(X,\mathcal{B},\mu,T)$ and any $A\in\mathcal{B}$ with $\mu(A)>0$ we have
@@
[…]

### v3

@@ 1.  Introduction
˜1.2 was later improved in [BLL08] to give an “if and only if” condition. A finite collection of polynomials $q_{1},\ldots,q_{k}\in\mathbb{Z}[t]$ is called jointly intersective if for all $m\in\mathbb{N}$ there is $n\in\mathbb{N}$ such that $q_{1}(n)\equiv\ldots\equiv q_{k}(n)\equiv{0}\bmod{m}$.
@@
In view of the above results, one is led to inquire whether similar results hold for more general classes of, say, eventually monotone sequences that do not grow too fast [^2] We remark in passing that for any sequence of exponential growth $q(n)$ there exists a set $E\subset\mathbb{N}$ with $\overline{\mathrm{d}}(E)>0$ which contains no pair of the form $\{x,x+q(n)\}$ with $x,n\in\mathbb{N}$ (cf. [BBHS08, Corollary 4.18]).. A natural class of sequences to consider are those arising from Hardy fields.
@@
Let $\mathsf{G}$ denote the ring (under pointwise addition and multiplication) of germs at infinity [^3] A germ at infinity is any equivalence class of real-valued functions in one real variable under the equivalence relationship $(f\equiv g)\Leftrightarrow\big(\exists t_{0}>0~\text{such that}~f(t)=g(t)~\text{for all}~t\in[t_{0},\infty)\big)$. of real valued functions defined on a half-line $s,\infty)$ for some $s\in\mathbb{R}$. Any subfield of $\mathsf{G}$ that is closed under differentiation is called a Hardy field.
By abuse of language, we say that a function $f\colon[s,\infty)\to\mathbb{R}$ belongs to some Hardy field $\mathcal{H}$, and write $f\in\mathcal{H}$, if its germ at infinity belongs to $\mathcal{H}$. For convenience, we assume all Hardy fields considered in this paper are shift-invariant, i.e., if $f\in\mathcal{H}$ then $f_{u}\in\mathcal{H}$ for all $u\in\mathbb{R}$, where $f_{u}(t)=f(t+u)$. A classical example of a (shift-invariant) Hardy field is the class of logarithmico-exponential functions introduced by Hardy in [Har12, Har71]. It consists of all (germs of) real-valued functions that can be built from real polynomials, the logarithmic function $\log(t)$, and the exponential function $\exp(t)$ using the standard arithmetical operations ++, -, $\cdot$, $\div$ and the operation of composition. Examples of logarithmico-exponential functions are p(t)/q(t){p(t)}/{q(t)} for $p(t),q(t)\in\mathbb{R}[t]$, $t^{c}$ for $c\in\mathbb{R}$, ${t}/{\log(t)}$ and $e^{\sqrt{t}}$, as well as any products or linear combinations of the above. Other Hardy fields contain even more exotic functions, such as $e^{\sqrt{\log t}}\Gamma(t)$ and $t^{c}\zeta(t)$ for any $c\in\mathbb{R}$, where $\zeta$ is the Riemann zeta function and $\Gamma$ is the usual gamma function (cf. [Bos84]). For more information on Hardy fields we refer the reader to [Bos81, Bos94, Fra09].
@@
Another variant of Szemerédi’s Theorem, which was recently obtained by the authors in [BMR20], stands in general position to ˜1.5 and reveals a new phenomenon pertaining to multiple recurrence along a wide family of non-polynomial (and not necessarily Hardy) functions. The following is one of the combinatorial corollaries of the main result in [BMR20].
@@
Both Theorems 1.5 and 1.6 leave some space for further inquiry. For instance, one would like to know if ˜1.6 has a version for several functions from $\mathcal{H}$ which deals with the patterns
@@
As for ˜1.5, a natural generalization is addressed by the following conjecture of Frantzikinakis. Given a finite collection of functions $f_{1},\ldots,f_{k}$, define
$\mathrm{span}(f_{1},\ldots,f_{k})\coloneqq\big\{c_{1}f_{1}(t)+\ldots+c_{k}f_{k}(t):(c_{1},\ldots,c_{k})\in\mathbb{R}^{k}\big\},$
@@
[…]
@@
[…]
The following theorem is the main combinatorial result of this paper. It contains Theorems 1.1, 1.2, 1.3, 1.5, 1.6 as special cases and confirms ˜1.7. We denote by $\mathrm{poly}(f_{1},\ldots,f_{k})$ the set of all real polynomials that can be “generated” using linear combinations of the functions $f_{1},\ldots,f_{k}$. More precisely,
@@
Condition (1) in Theorem A will still hold if all the $f_{i}$ are shifted. Together with ˜1.8, this observation implies that if we are in case (1) of Theorem A then the conclusion of Theorem A remains true, even if the closest integer function $\cdot$] is replaced by either $\lfloor.\rfloor$ or $\lceil.\rceil$ (or indeed with any other rounding function of the form $\lfloor x+c\rfloor$, $c\in[0,1)$). This observation, however, does not apply to condition (2) – see ˜1.14 below.
Besides resolving ˜1.7, Theorem A also implies numerous new results. Corollaries A1, A2, A3, and A4 below comprise a selection of such results that we consider to be of particular interest.
The following rather special case of Theorem A already implies Theorems 1.1, 1.2, 1.5, and 1.6.
@@
where $c_{1},\ldots,c_{k}$ are all positive integers, follows from ˜1.2, whereas the case when $c_{1},\ldots,c_{k}$ are all positive non-integers follows from ˜1.5. The next corollary of Theorem A deals with the previously unknown case where the constants $c_{1},\ldots,c_{k}$ are a mix of integers and non-integers.
@@
For any $c_{1},\ldots,c_{k}>0$ and any set $E\subset\mathbb{N}$ of positive upper density there exist $a,n\in\mathbb{N}$ such that $\{a,\,a+[n^{c_{1}}],\ldots,a+[n^{c_{k}}]\}\subset E$. The same is true with $[.]$ replaced by either $\lfloor.\rfloor$ or $\lceil.\rceil$.
@@
We remark that Corollaries A1 and A3 are not true if $[.]$ is replaced by either $\lfloor.\rfloor$ or $\lceil.\rceil$ (see ˜1.14 below).
While Corollary A3 contains ˜1.2 as a special case, it does not encompass polynomials which have a non-zero constant term. The following theorem shows that the conclusion of Corollary A3 holds for significantly more general families $\mathcal{K}$}. In particular, in conjunction with Theorem A, it implies both ˜1.3 and ˜1.5.
@@
Let $q\in\mathbb{Z}[t]$ be an intersective polynomial, let $\mathcal{H}$ be a Hardy field and let $\mathcal{L}$$\subset\mathcal{H}$ be a family of functions such that any $f\in{\mathcal{L}}$ satisfies $t^{k-1}\prec f(t)\prec t^{k}$ for some $k\in\mathbb{N}$, and any distinct $f,g\in{\mathcal{L}}$ have different growth. Let ${\mathcal{K}}({\mathcal{L}},q)$ be the linear span over $\mathbb{R}$ of $\mathcal{L}$} and $q\mathbb{R}[t]$. Then any tuple $f_{1},\dots,f_{k}$ from ${\mathcal{K}}({\mathcal{L}},q)$ satisfies Condition (2) of Theorem A.
@@
-$\nabla\text{-}\mathrm{span}(f_{1},\ldots,f_{k})$ $\coloneqq\mathrm{span}\big(\{f_{i}^{(m)}:1\leqslant i\leqslant k;\ m\geqslant 0\}\big)$
Theorem A allows us to derive a corollary which extends ˜1.6 from a single function $f$ to multiple functions $f_{1},\ldots,f_{k}$.
@@
The same is true with $[.]$ replaced by either $\lfloor.\rfloor$ or $\lceil.\rceil$.
The assumptions of Corollary A4 are satisfied, for instance, if all the $f_{i}$ are linear combinations of powers $t^{c}$ with non-integer exponents $c>0$. On the other hand, if some $f\in\mathrm{span}^{*}(f_{1},\dots,f_{k})$ is a polynomial, then the conclusion of Corollary A4 fails. In Example 1.20 below we show that the conclusion may fail even when $\mathrm{poly}(f_{1},\dots,f_{k})=\emptyset$.
@@ 1.2.  Ergodic results
In light of (1.2), it is natural to tackle Theorem A by studying the behaviour of Cesàro averages of multicorrelation expressions of the form
@@ 1. Property (P):
It is important to mention that given a finite collection of functions $f_{1},\ldots,f_{k}\in\mathcal{H}$ of polynomial growth from a Hardy field $\mathcal{H}$, not every $W\in\mathcal{H}$ will satisfy property (P). For example $W(t)=t$ is not compatible with $f(t)=\log t$. On the other hand, there always exists a compatible $W\in\mathcal{H}$ (see ˜2.2). Roughly speaking, $W$ can be chosen as any function which tends to $\infty$ and grows “slower” than any unbounded function in
@@
Let $f_{1},\ldots,f_{k}$ be functions of polynomial growth from a Hardy field $\mathcal{H}$ and let $W\in\mathcal{H}$ be any function with $1\prec W(t)\ll t$ and such that $f_{1},\ldots,f_{k}$ satisfy Property (P). Let $(X,\mathcal{B},\mu,T)$ be an invertible measure preserving system.
  1. (i)
@@
  2. (ii)
    If $f_{1},\ldots,f_{k}$ satisfy condition (1) of Theorem A then for any $h_{1},\ldots,h_{k}\in L^{\infty}(X)$
    $\lim_{N\to\infty}\frac{1}{W(N)}\sum_{n=1}^{N}w(n)T^{[f_{1}(n)]}h_{1}\cdots T^{[f_{k}(n)]}h_{k}\,=\,\prod_{i=1}^{k}h_{i}^{*}\quad\text{in}~L^{2},$ (1.6)
@@
  3. (iii)
    If $f_{1},\ldots,f_{k}$ satisfy condition (2) of Theorem A then for any $A\in\mathcal{B}$ with $\mu(A)>0$ we have
    $\lim_{N\to\infty}\frac{1}{W(N)}\sum_{n=1}^{N}w(n)\mu\Big(A\cap T^{-[f_{1}(n)]}A\cap\ldots\cap T^{-[f_{k}(n)]}A\Big)\,>\,0.$ (1.7)
As a first corollary to Theorem B we obtain convergence of multiple ergodic averages with Cesàro weights along a rather large class of functions from a Hardy field. Special cases of this corollary were previously obtained in [HK05b, HK05a, Lei05, BHK09, Fra10, Kou18].
@@
exists in $L^{2}$. The same is true with $[.]$ replaced by either $\lfloor.\rfloor$ or $\lceil.\rceil$.
Theorem B also provides additional information on the following result, which was originally conjectured by Frantzikinakis ([Fra16, Problem 23]; cf. also [Fra15, Problem 1]) and recently proved by Tsinas [Tsi23].
@@
Part (ii) of Theorem B implies ˜1.12 under the additional assumption that the functions $f_{1},\dots,f_{k}$ satisfy Property (P) with $W(t)=t$, which is slightly stronger than the assumption that $|f(t)-q(t)|/\log t\to\infty$ for every $q\in\mathbb{Z}[t]$ and $f\in\mathrm{span}^{*}(f_{1},\ldots,f_{k})$. However, Theorem B gives more than that. Indeed, part (ii) of Theorem B provides a confirmation of a variant of ˜1.12 where the condition $|f(t)-q(t)|/\log t\to\infty$ is weakened to $|f(t)-q(t)|\to\infty$, but at the price of replacing Cesàro averages by weighted averages (where one can choose any weight for which Property (P) is satisfied, cf. ˜2.2).
@@
Let $f_{1},\ldots,f_{k}$ be functions of polynomial growth from a Hardy field $\mathcal{H}$ and let $W\in\mathcal{H}$ be any function with $1\prec W(t)\ll t$ and such that $f_{1},\ldots,f_{k}$ satisfy Property (P). Suppose that for all $q\in\mathbb{Z}[t]$ and $f\in\mathrm{span}^{*}(f_{1},\ldots,f_{k})$ we have $\lim_{t\to\infty}|f(t)-q(t)|=\infty$. Then for any invertible ergodic measure preserving system $(X,\mathcal{B},\mu,T)$ and any $h_{1},\dots,h_{k}\in L^{\infty}(X)$
@@
The same is true when $[.]$ is replaced by either $\lfloor.\rfloor$ or $\lceil.\rceil$.
@@
Finally, we formulate two corollaries of Theorem B which imply Theorem A (via Furstenberg’s correspondence principle). We believe that these two results are also of independent interest.
@@
Let $f_{1},\ldots,f_{k}$ be functions of polynomial growth from a Hardy field $\mathcal{H}$ satisfying condition (1) of Theorem A. Then for any invertible measure preserving system $(X,\mathcal{B},\mu,T)$ and any $A\in\mathcal{B}$,
@@
The same is true when $[.]$ is replaced by either $\lfloor.\rfloor$ or $\lceil.\rceil$.
@@
Let $f_{1},\ldots,f_{k}$ be functions of polynomial growth from a Hardy field $\mathcal{H}$ satisfying condition (2) of Theorem A. Then for any invertible measure preserving system $(X,\mathcal{B},\mu,T)$ and any $A\in\mathcal{B}$ with $\mu(A)>0$ we have
@@
We would like to stress that one cannot replace $\limsup$ with $\lim$ in (1.9) and (1.10), as the limits don’t exist in general. In fact, if one replaces $\limsup$ with $\liminf$, the left hand side may equal 0 in either case.